OpenStudy (yato_gami):

I need help with this please! I want someone to go through the steps with me, not tell me the answer. Thanks~! ^_^ The amount of a radioactive material changes with time. The table below shows the amount of radioactive material f(t) left after time t: t(hours) 0 1 2 f(t) 100 50 25 Which exponential function best represents the relationship between f(t) and t?

OpenStudy (yato_gami):

f(t) = 0.5(100)t f(t) = 100 (0.25)t f(t) = 0.25(50)t f(t) = 100(0.5)t

OpenStudy (yato_gami):

@Directrix Can you help with this?

OpenStudy (welshfella):

Let f(t) = a x^t first plug in t = 0 and f(t) = 100 and find a . What do you get?

OpenStudy (welshfella):

100 = a x^0 right?

OpenStudy (welshfella):

- because when t = 0 , f(t) = 100

OpenStudy (yato_gami):

Alright I got it so far.

OpenStudy (yato_gami):

Is that it?